OpenStudy (saad560):

3+q+9=12 q=18 q=0 q=6 q=24

OpenStudy (saad560):

@triciaal

OpenStudy (awesomedude2078):

3+9=12, so q+12=12

alones (alones):

So you'll have to start from simplify both sides of the equation \(\ 3+q+9=12\) \(\ q+12=12\) than subtract \(\ 12\) from both sides \(\ q+12−12=12−12\) So you'll end up with \(\ q=0\)